# On the applicability of Kolmogorov’s theory of probability

### v2

@@ 3 Physical considerations
Second, the most direct result on the applicability of Kolmogorov’s axioms to quantum phenomena, which I am aware of, is the BCHSH inequality [clauserProposedExperimentTest1969, morettiFundamentalMathematicalStructures2019]. This constitutes a generalization of Bell’s theorem [bellEinsteinPodolskyRosen1964], that overcomes the “unrealistic restriction that for some pair of parameters […] there is perfect correlation” [clauserProposedExperimentTest1969]. If one follows the statement that those inequalities are empirically violated, then, by symmetry considerations, for any unit vectors $\vec{a}$, $\vec{b}\in\mathbb{R}^{3}$ the random variable $X(\vec{a},\vec{b},\,.\,)$ in question must be
$X(\vec{a},\vec{b},\,.\,)=A(\vec{a},\vec{b},\,.\,)\,B(\vec{a},\vec{b},\,.\,)$ (3.2)
instead of $A(\vec{a},\,.\,)\,B(\vec{b},\,.\,)$. Whether the approach described here supports this kind of “(Bell) non-locality” needs to be investigated in the context of a respective dynamical theory [takabayasiDescriptionPauliMatter1957, janossyHydrodynamicalModelWave1966]. $\diamondsuit$
@@ 4 Born’s rule as an approach for basing quantum theory on Kolmogorov’s axioms
The Born rule for position provides a natural approach to the construction of a quantum theory on the basis of Kolmogorov’s axioms. [^8] See [reddigerProbabilisticFoundationRelativistic2024] for how this is to be done in the (general-)relativistic domain. In what follows I shall show how this is to be done for an $N$-body quantum system ($N\in\mathbb{N}$).
Consider a normalized (Schrödinger) wave function $\Psi_{0}\in L^{2}(\mathbb{R}^{3N},\mathbb{C})$ on the configuration space $\mathbb{R}^{3N}$ and denote by $\mathcal{B}^{*}(\mathbb{R}^{3N})$ the corresponding $\sigma$-algebra of Lebesgue sets. Then, by Born’s rule, the probability that the bodies are in a configuration lying in $W\in\mathcal{B}^{*}(\mathbb{R}^{3N})$ is
$\mathbb{P}_{0}\left(W\right)=\int_{W}\left\lvert\Psi_{0}\right\rvert^{2}\negthinspace\left(\vec{r}_{1},\dots,\vec{r}_{N}\right)\,\mathrm{d}^{3N}r\ .$ (4.1)
Indeed, normalization of $\Psi_{0}$ is equivalent to $\mathbb{P}_{0}(\mathbb{R}^{3N})=1$. Furthermore, if $t\mapsto U_{t}$ is any $1$-parameter family of unitary evolution operators determining the time evolution of $\Psi_{0}$ via $\Psi_{t}=U_{t}\Psi_{0}$ for $t\in\mathbb{R}$, we can use $\Psi_{t}$ to define $\mathbb{P}_{t}$ in analogy to Eq. (4.1). Then unitarity of $U_{t}$ implies that $\mathbb{P}_{t}(\mathbb{R}^{3N})=1$ for all $t\in\mathbb{R}$. Accordingly, the following holds.
@@
The tuple $\bigl{(}\mathbb{R}^{3N},\mathcal{B}^{*}(\mathbb{R}^{3N}),\mathbb{P}_{t}\bigr{)}$ is a probability space for every $t\in\mathbb{R}$. $\diamondsuit$
From a mathematical point of view, Lem. 1 is all one requires to base any such $N$-body quantum theory on the basis of Kolmogorov’s axioms. By applying standard probability theory [klenkeProbabilityTheoryComprehensive2013], Lem. 1 determines what a random variable is, how their expectation values are defined, how conditional probabilities are obtained, etc.
@@
The dynamical evolution $t\mapsto U_{t}$ one can generally carry over from QM. The reason is, that this is less difficult to test empirically than the underlying probability theory: a physically incorrect evolution tends to predict physically incorrect emission spectra, scattering cross sections etc., so that in this respect quantum-mechanical models are comparatively trustworthy.
The actual difficulties concern the observables and the status of the projection postulate. In particular, the ‘observables’ ought to be real random variables on $\bigl{(}\mathbb{R}^{3N},\mathcal{B}^{*}(\mathbb{R}^{3N})\bigr{)}$ instead of (unbounded, densely defined) self-adjoint operators on $L^{2}(\mathbb{R}^{3N},\mathbb{C}^{k})$. In what follows I shall motivate and rigorously define such random variables on the basis of physical considerations. The projection postulate and the question of measurement shall be addressed via conditional probabilities in Part II of this series.
@@ 5 Madelung’s equations
In 1926 Madelung discovered a reformulation of the $1$-body Schrödinger equation, which is reminiscent Newtonian continuum mechanics [madelungAnschaulicheDeutungGleichung1926, madelungQuantentheorieHydrodynamischerForm1927, nelsonDerivationSchrodingerEquation1966, vonrenesseOptimalTransportView2012, reddigerMadelungPictureFoundation2017]. His reformulation can be generalized to the $N$-body Schrödinger equation, which is what I shall consider here (cf. [janossyHydrodynamicalModelWave1969] and §7.1.2 in [hollandQuantumTheoryMotion1993]).
Assuming that the time-dependent wave function $\Psi$ is smooth, we set
@@
Those $N$-body Madelung equations, Eqs. (5.4), are dynamical equations for the time-dependent particle position probability density $\rho$ and the so called drift (velocity) field $\vec{v}=(\vec{v}_{1},\dots,\vec{v}_{N})$. In particular, the continuity equation, Eq. (5.4b), assures probability conservation and states that the velocity vector field $\vec{v}$ governs the evolution of particle position probability for the ensemble (cf. Sec. 5.1 of [reddigerMadelungPictureFoundation2017]). Eq. (5.4a) is naturally understood as a force equation for the $a$th body and Eqs. (5.4c) constitute a condition of vanishing ‘quantum vorticity’ [reddigerMathematicalTheoryMadelung2023]. The Madelung equations are also employed in de Broglie-Bohm theory [bohmSuggestedInterpretationQuantum1952, bohmSuggestedInterpretationQuantum1952a, hollandQuantumTheoryMotion1993, bohmUndividedUniverseOntological1993] and stochastic mechanics [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952, nelsonDerivationSchrodingerEquation1966, nelsonQuantumFluctuations1985].
The overt analogy to Newtonian continuum mechanics – more specifically, the theory of Brownian motion, as first observed by Fényes [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952] – provides us with a number of physically natural random variables.
In this $N$-body theory they are the position of the $a$th body $\vec{r}_{a}$, its (drift) momentum $\vec{p}_{a}=m_{a}\vec{v}_{a}$, its angular momentum $\vec{l}_{a}=(\vec{r}_{a}-\vec{r}_{0})\times\vec{p}_{a}$ around a given $\vec{r}_{0}\in\mathbb{R}^{3}$, and the total energy, as obtained from Eq. (5.4a): [^9] In addition, it is worth considering the stochastic/osmotic velocity for the $a$th body $\vec{u}_{a}=\frac{\hbar}{2m_{a}}\nabla_{a}\rho/\rho$. See e.g. Refs. [nelsonDerivationSchrodingerEquation1966, delapenaDoesQuantumMechanics1982, henkMachineLearningQuantum2023, reddigerMathematicalTheoryMadelung2023].
@@
Notable derived quantities are the center of mass, the total momentum, and the total angular momentum (around any $\vec{r}_{0}$).
@@
Following this line of reasoning, it may seem like a ‘miracle’ that the following holds: The (Kolmogorovian) expectation value of any one of the aforementioned random variables is exactly equal to the expectation value of the corresponding quantum-mechanical operator. For the $1$-body theory this was first observed by Takabayasi [takabayasiFormulationQuantumMechanics1952]. My contribution to this debate is the recognition that this constitutes evidence for the validity of Kolmogorov’s theory of probability in the quantum domain (cf. [reddigerMadelungPictureFoundation2017] and Sec. 2.3 in [reddigerProbabilisticFoundationNonrelativistic2022]).
@@
In this approach ‘quantization’ still occurs: if one of the aforementioned random variables is a constant, then the induced probability measure is a Dirac measure. This is the case, for instance, for $E$ from Eq. (5.5), whenever $\Psi$ is a stationary state in the quantum-mechanical sense.
@@
Of course, to obtain a mathematically rigorous theory a restriction to local statements and requiring smoothness is not enough. Furthermore, if one ascribes such physical significance to the Madelung equations, then it is natural to build the theory thereon [reddigerMadelungPictureFoundation2017]. To this end, in [reddigerMathematicalTheoryMadelung2023] the mathematical theory of the Madelung equations was reviewed and it was concluded that this theory is still underdeveloped. While Gasser and Markowich did provide a suitable weak formulation [gasserQuantumHydrodynamicsWigner1997], it is not entirely understood, how the Schrödinger equation can be reobtained (for $H^{1}$-wave functions) [reddigerMathematicalTheoryMadelung2023]. [^10] This would also resolve the question of the “classical limit” in a natural manner, see [gasserQuantumHydrodynamicsWigner1997], Rem. 3.1 in [reddigerMathematicalTheoryMadelung2023], and Sec. 5.1 in [reddigerMadelungPictureFoundation2017].
@@
Fortunately, for the purpose of constructing a mathematically rigorous quantum theory on the basis of Kolmogorov’s axioms this does not pose a direct problem. Taking a pragmatic approach, we can carry over the dynamical evolution $t\mapsto U_{t}$ from QM and then only the corresponding physical random variables need to be rigorously defined. This may be understood as a hybrid theory, for it still borrows concepts from QM—as opposed to erecting the entire theory on Madelung’s equations and the Born rule in conjunction with Kolmogorov’s axioms [reddigerMadelungPictureFoundation2017].
@@ 6 A hybrid theory
For the purpose of formulating the hybrid theory, we shall constrain ourselves to (time-independent) Hamiltonians of the form
@@
with $V$ real-valued and Lebesgue measurable. We further require that $\hat{H}$ is self-adjoint on some dense domain $\operatorname{dom}\hat{H}$ in $L^{2}(\mathbb{R}^{3N},\mathbb{C})$ and that $\operatorname{dom}\hat{H}\subseteq H^{2}(\mathbb{R}^{3N},\mathbb{C})$.
@@
The above restriction is made, because ‘observables’ always relate to particular models and it is the ‘observables’ relating to (the weak formulation of) Eq. (5.3) we consider here. Nonetheless, the main result in this context, Thm. 1 below, can in principle be generalized to other classes of models, though ultimately physical considerations dictate what random variables one ought to consider. [^11] The definition of the energy in Eq. (6.3c) below is model-independent and an analogous definition may be used for time-dependent Hamiltonians (see e.g. [schmidKatoTheoremIntegration2014], Sec. V in [penzRegularityEvolutionEquations2018], and [balmasedaSchrodingerEquationTimeDependent2022]).
@@
The reader may have noticed that QM makes weaker assumptions on the regularity of wave functions. However, for $L^{2}$-wave functions the existence of the expectation value of e.g. the quantum-mechanical energy is generally not guaranteed, so that even in QM some assumptions of weak differentiability are physically justifiable (cf. Refs. [grublNondifferentiableBohmianTrajectories2011, berryQuantumFractalsBoxes1996, reddigerMathematicalTheoryMadelung2023, carlenConservativeDiffusions1984]). Still, it may be possible to generalize Thm. 1, so that the required order of differentiability on $\Psi_{0}$ is reduced by $1$ (cf. Appx. A in [reddigerMathematicalTheoryMadelung2023]).
Similarly, the fact that $\mathbb{E}_{t}(\vec{r}_{a})$ or $\mathbb{E}_{t}(\vec{l}_{a})$ need not be defined for a given $\Psi_{t}$ does not pose any conceptual problems: they are defined if and only if the analogous quantum-mechanical expectation values are—the respective quantities are equal.
@@
We consider the simplest quantum-mechanical model of the hydrogen atom, i.e. $N=1$ and $V(\vec{r})=-\text{e}^{2}/(4\pi\varepsilon_{0}\left\lvert\vec{r}\right\rvert)$. Using physics conventions, we denote the electron mass by $\mu$ and each stationary state basis function by $\Psi_{nlm}$; $n$, $l$, and $m$ are appropriately chosen integers.
For any such integers we indeed have $\Psi_{nlm}\in\operatorname{dom}\hat{H}=H^{2}(\mathbb{R}^{3},\mathbb{C})$. Recalling that $\Psi_{nlm}(t,\vec{r})\equiv e^{-\mathfrak{i}E_{n}t/\hbar}\,\Psi_{nlm}(\vec{r})$, the respective random variables from Thm. 1.2 are given via
@@
Finally, I would like to remark that the ‘observables’ $\vec{v}_{a}$ and $E$ from Thm. 1 depend on the particular ensemble one considers (as encoded by $\Psi_{t}$). This is in stark contrast to the momentum and energy operator in QM, which generally only depend on the model, not the particular ensemble.

### v5

@@ 3 Physical considerations
Second, the most direct result on the applicability of Kolmogorov’s axioms to quantum phenomena, which the author is aware of and which is empirically relevant, is the BCHSH inequality (cf. [clauserProposedExperimentTest1969, morettiFundamentalMathematicalStructures2019] and also Sec. 8 in [gillStatisticsCausalityBells2014] for a refutation of GHZ-type arguments). This constitutes a generalization of Bell’s theorem [bellEinsteinPodolskyRosen1964], that overcomes the “unrealistic restriction that for some pair of parameters […] there is perfect correlation” [clauserProposedExperimentTest1969]. The latter can be derived on the basis of Kolmogorov’s axioms and the additional assumption of Bell locality: each detector random variable $A$ and $B$ only depends on the setting of the respective detector, not on the setting of the other.
Though there are several variations of such Bell inequalities, it is generally acknowledged that they are empirically violated [aspectExperimentalTestsRealistic1981, aspectExperimentalTestBell1982, aspectClosingDoorEinstein2015, matsukevichBellInequalityViolation2008, hensenLoopholefreeBellInequality2015, rosenfeldEventReadyBellTest2017]. It is noteworthy that such results are necessarily of statistical nature and that statistical means of sample data generated by “local realistic” models need not always satisfy a respective inequality for the means. Still, provided that the trials are mutually independent and the products of $A$ and $B$ are identically distributed for each setting, the central limit theorem makes the empirical investigation of the violation of Bell inequalities feasible. Indeed, some experiments were able to meet the conventional threshold of a violation by $5$ standard deviations.
Part II of this series reviews the underlying mathematical argument and proposes that Bell nonlocality may be a genuine physical phenomenon, resulting out of quantum dynamics. $\diamondsuit$
@@ 4 Born’s rule as an approach for basing quantum theory on Kolmogorov’s axioms
The Born rule for position provides a natural approach to the construction of a quantum theory on the basis of Kolmogorov’s axioms [reddigerProbabilisticFoundationRelativistic2024]. This section shows how to do this for an $N$-body quantum system “without spin” ‘($N\in\mathbb{N}$).
We consider a time-dependent, normalized (Schrödinger) wave function $\Psi_{t}\in L^{2}(\mathbb{R}^{3N},\mathbb{C})$ on the configuration space $\mathbb{R}^{3N}$ and denote by $\mathcal{B}^{*}(\mathbb{R}^{3N})$ the corresponding $\sigma$-algebra of Lebesgue sets. Then, by Born’s rule, the probability that the bodies are in a configuration lying in $W\in\mathcal{B}^{*}(\mathbb{R}^{3N})$ at time $t\in\mathbb{R}$ is
$\mathbb{P}_{t}\left(W\right)=\int_{W}\left\lvert\Psi_{t}\right\rvert^{2}\negthinspace\left(\vec{r}_{1},\dots,\vec{r}_{N}\right)\,\mathrm{d}^{3N}r\ .$ (4.1)
Indeed, normalization of $\Psi_{t}$ is equivalent to $\mathbb{P}_{t}(\mathbb{R}^{3N})=1$. Normalization, in turn, follows from unitary evolution: If, initially, $\Psi_{0}\in L^{2}(\mathbb{R}^{3N},\mathbb{C})$ and $t\mapsto U_{t}$ is any $1$-parameter family of unitary evolution operators, then $\Psi_{t}=U_{t}\Psi_{0}$ is indeed normalized for any $t\in\mathbb{R}$. Accordingly, the following holds.
@@
Let $N\in\mathbb{N}$, $\Psi_{0}\in L^{2}(\mathbb{R}^{3N},\mathbb{C})$, and for all $t\in\mathbb{R}$ let $U_{t}$ be a unitary operator on $L^{2}(\mathbb{R}^{3N},\mathbb{C})$. Set $\Psi_{t}\equiv U_{t}\Psi_{0}$ and define the measure $\mathbb{P}_{t}$ on $\mathcal{B}^{*}(\mathbb{R}^{3N})$ via Eq. (4.1).
Then the tuple $\bigl{(}\mathbb{R}^{3N},\mathcal{B}^{*}(\mathbb{R}^{3N}),\mathbb{P}_{t}\bigr{)}$ is a probability space for every $t\in\mathbb{R}$. $\diamondsuit$
From a mathematical point of view, Lem. 1 is all we require to base any such $N$-body quantum theory on the basis of Kolmogorov’s axioms. By applying standard probability theory [klenkeProbabilityTheoryComprehensive2013], Lem. 1 determines what a random variable is, how their expectation values and distributions are defined, how conditional probabilities are obtained, etc.
@@
The dynamical evolution $t\mapsto U_{t}$ one can generally carry over from quantum mechanics. The reason is, that this is less difficult to test empirically than the underlying probability theory: a physically incorrect evolution tends to predict physically incorrect emission spectra, scattering cross sections etc., so that, in this respect, quantum-mechanical models are comparatively trustworthy.
The actual difficulties concern the observables and the status of the projection postulate. In particular, the ‘observables’ ought to be real random variables on $\bigl{(}\mathbb{R}^{3N},\mathcal{B}^{*}(\mathbb{R}^{3N})\bigr{)}$ instead of (unbounded, densely defined) self-adjoint operators on $L^{2}(\mathbb{R}^{3N},\mathbb{C})$. In the next section we motivate such random variables on the basis of physical considerations. In Sec. 6 thereafter, those random variables are defined rigorously. The projection postulate and the question of measurement are addressed via conditional probabilities in Part III of this series.
@@ 5 Madelung’s equations
In 1926 Madelung discovered a reformulation of the $1$-body Schrödinger equation, which is reminiscent Newtonian continuum mechanics [madelungAnschaulicheDeutungGleichung1926, madelungQuantentheorieHydrodynamischerForm1927, nelsonDerivationSchrodingerEquation1966, vonrenesseOptimalTransportView2012, mitaSchrodingersEquationDiffusion2021, reddigerMadelungPictureFoundation2017]. His reformulation can be generalized to the $N$-body Schrödinger equation, which is what I shall consider here (cf. [janossyHydrodynamicalModelWave1969] and §7.1.2 in [hollandQuantumTheoryMotion1993]).
Assuming, for the sake of simplicity, that the time-dependent wave function $\Psi$ is smooth, we set
@@
Those $N$-body Madelung equations, Eqs. (5.4), are dynamical equations for the time-dependent particle position probability density $\rho$ and the (probability) current velocity (vector field) $\vec{v}=(\vec{v}_{1},\dots,\vec{v}_{N})$ (following Nelson’s terminology [nelsonDerivationSchrodingerEquation1966]).
In particular, the continuity equation, Eq. (5.4b), assures probability conservation and states that the current velocity $\vec{v}$ governs the evolution of particle position probability for the ensemble (cf. Sec. 5.1 in [reddigerMadelungPictureFoundation2017]). Eq. (5.4a) is naturally understood as a force equation for the $a$th body and Eqs. (5.4c) constitute a condition of vanishing ‘quantum vorticity’ [reddigerMathematicalTheoryMadelung2023]. The Madelung equations are also employed in de Broglie-Bohm theory [bohmSuggestedInterpretationQuantum1952, bohmSuggestedInterpretationQuantum1952a, hollandQuantumTheoryMotion1993, bohmUndividedUniverseOntological1993] and stochastic mechanics [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952, nelsonDerivationSchrodingerEquation1966, nelsonQuantumFluctuations1985].
There is an overt analogy to Newtonian continuum mechanics or, more specifically and as first observed by Fényes [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952], the theory of Brownian motion. This analogy provides us with a number of physically natural random variables.
In this $N$-body theory they are the position of the $a$th body $\vec{r}_{a}$, its (current) momentum $\vec{p}_{a}=m_{a}\vec{v}_{a}$, its angular momentum $\vec{l}_{a}=(\vec{r}_{a}-\vec{r}_{0})\times\vec{p}_{a}$ around a given $\vec{r}_{0}\in\mathbb{R}^{3}$, and the total energy, as obtained from Eq. (5.4a):
@@
In addition, one may introduce the stochastic velocity (vector field) for the $a$th body
$\vec{u}_{a}=\frac{\hbar}{m_{a}}\operatorname{Re}(\nabla_{a}\Psi/\Psi)=\frac{\hbar}{2m_{a}}\nabla_{a}\rho/\rho\ .$ (5.6)
The physical relevance of $\vec{u}=(\vec{u}_{1},\dots,\vec{u}_{N})$ is justified on the grounds, that the so called quantum force [hollandQuantumTheoryMotion1993] on the right hand side of Eq. (5.4a) may be reformulated in terms of it and that the expectation value of the quantum-mechanical kinetic energy contains a summand of the type $\sum_{a=1}^{N}\frac{m_{a}}{2}\,\vec{u}_{a}^{2}$ (see for example [nelsonDerivationSchrodingerEquation1966, delapenaDoesQuantumMechanics1982, henkMachineLearningQuantum2023, carlenConservativeDiffusions1984, reddigerMathematicalTheoryMadelung2023]).
@@
Notable derived quantities are the center of mass, the total momentum, and the total angular momentum (around any $\vec{r}_{0}$). How one defines kinetic energies ultimately depends on the particular interpretation of the stochastic velocity and we shall not commit to any particular one here.
@@
Following this line of reasoning and putting the topic of the stochastic velocity aside for a moment, it may seem like a ‘miracle’ that the following holds: The (Kolmogorovian) expectation value of any one of the aforementioned random variables is exactly equal to the expectation value of the corresponding quantum-mechanical operator. For the $1$-body theory this was first observed by Takabayasi [takabayasiFormulationQuantumMechanics1952]. The author’s contribution to this debate is the recognition that this constitutes evidence for the validity of Kolmogorov’s theory of probability in the quantum domain (cf. [reddigerMadelungPictureFoundation2017] and Sec. 2.3 in [reddigerProbabilisticFoundationNonrelativistic2022]).
@@
In this approach ‘quantization’ still occurs: if one of the aforementioned random variables is a constant, then the induced probability measure is a Dirac measure. This is the case, for instance, for $E$ from Eq. (5.5), whenever $\Psi$ is a stationary state in the quantum-mechanical sense. As shown in Sec. 6 below, the notion of a stationary wave function may also be justified within this approach.
@@
Of course, to obtain a mathematically rigorous theory, a restriction to local statements and requiring smoothness is not enough. Furthermore, if one ascribes such physical significance to the Madelung equations, then it is natural to build the theory thereon. In [reddigerMadelungPictureFoundation2017] this theory was named geometric quantum theory. To this end, R. and Poirier reviewed the mathematical theory of the Madelung equations in [reddigerMathematicalTheoryMadelung2023], concluding that the mathematical theory of the Madelung equations is still underdeveloped. While Gasser and Markowich did provide a suitable weak formulation [gasserQuantumHydrodynamicsWigner1997], it is not entirely understood, how the Schrödinger equation can be reobtained [reddigerMathematicalTheoryMadelung2023]. As a side note, the work by Gasser and Markowich is of particular interest in this context, since the authors showed how to resolve the question of the “classical limit” in a natural manner (see [gasserQuantumHydrodynamicsWigner1997], Rem. 3.1 in [reddigerMathematicalTheoryMadelung2023], and Sec. 5.1 in [reddigerMadelungPictureFoundation2017]).
@@
Fortunately, for the purpose of constructing a mathematically rigorous quantum theory on the basis of Kolmogorov’s axioms, the contemporary lack of a mathematical theory of the Madelung equations does not pose a direct problem. Taking a pragmatic approach, we can carry over the dynamical evolution $t\mapsto U_{t}$ from quantum mechanics and then only the corresponding physical random variables need to be rigorously defined. This may be understood as a hybrid theory, for it needs to rely on wave functions satisfying the Schrödinger equation as primary mathematical objects. This contrasts with the aim of geometric quantum theory as a research program, which is to erect the entire theory on Madelung’s equations and the Born rule in conjunction with Kolmogorov’s axioms [reddigerMadelungPictureFoundation2017].
@@ 6 A hybrid theory
For the purpose of formulating the hybrid theory, we constrain ourselves to (time-independent) Hamiltonians of the form
@@
with $V$ real-valued and Lebesgue measurable. We further require that $\hat{H}$ is self-adjoint on some dense domain $\operatorname{dom}\hat{H}$ in $L^{2}(\mathbb{R}^{3N},\mathbb{C})$ and that $\operatorname{dom}\hat{H}$ is a subset of the Sobolev space $H^{2}(\mathbb{R}^{3N},\mathbb{C})$ (cf. Supplement 2 in [berezinSchrodingerEquation1991]).
@@
The above restriction is made, because ‘observables’ always relate to particular models and it is the ‘observables’ relating to Eq. (5.3) (understood in a mathematically suitable sense) we consider here. Nonetheless, the main result in this context, Thm. 1 below, can in principle be generalized to other classes of models, though ultimately physical considerations dictate what random variables one ought to consider.
@@
Remark 2
  1. 1)
    In full analogy to the definition of $\vec{v}_{a}$ in Thm. 1.2, for $t\in\mathbb{R}$ we may define
    $(\vec{u}_{a})_{t}=\frac{\frac{\hbar}{m_{a}}\operatorname{Re}\left(\Psi_{t}^{*}\nabla_{a}\Psi_{t}\right)\mathrm{d}^{3N}r}{\mathrm{d}\mathbb{P}_{t}}\ .$ (6.4)
  2. 2)
    The definition of the energy in Eq. (6.3c) above is model-independent in the sense that it holds for any time-independent, densely defined, self-adjoint Hamiltonian $\hat{H}$ on $L^{2}(\mathbb{R}^{3N},\mathbb{C})$ and wave function $\Psi\colon t\mapsto\Psi_{t}$ satisfying Eq. (6.3a) with $\Psi_{0}\in\operatorname{dom}\hat{H}$. An analogous definition may be used for time-dependent Hamiltonians, though the treatment of domains requires more care (see for instance [schmidKatoTheoremIntegration2014], Sec. V in [penzRegularityEvolutionEquations2018], and [balmasedaSchrodingerEquationTimeDependent2022]).
$\diamondsuit$
The reader may have noticed that quantum mechanics makes weaker assumptions on the regularity of wave functions. However, for $L^{2}$-wave functions the existence of the expectation value of, for instance, the quantum-mechanical energy is generally not guaranteed, so that even in quantum mechanics some assumptions of weak differentiability are physically justifiable (cf. Refs. [grublNondifferentiableBohmianTrajectories2011, berryQuantumFractalsBoxes1996, reddigerMathematicalTheoryMadelung2023, carlenConservativeDiffusions1984]). Still, it may be possible to generalize Thm. 1, so that the required order of differentiability on $\Psi_{0}$ is reduced by $1$ (cf. Appx. A in [reddigerMathematicalTheoryMadelung2023]).
Similarly, the fact that $\mathbb{E}_{t}(\vec{r}_{a})$ or $\mathbb{E}_{t}(\vec{l}_{a})$ need not be defined for a given $\Psi_{t}$ does not pose any conceptual problems: they are defined if and only if the analogous quantum-mechanical expectation values are: the respective quantities are equal.
Before considering an explicit example, it is worth pointing out that both the time-dependent and the time-independent Schrödinger equation can be derived from the time evolution law in Eq. (6.3a). Thus, the theory predicts the same energies and stationary wave functions as quantum mechanics. Furthermore, in geometric quantum theory the notion of a stationary wave function may be justified physically by imposing the requirement that the family of probability measures $(\mathbb{P}_{t})_{t\in\mathbb{R}}$ is stationary. By definition, this holds, if $\mathbb{P}_{t}=\mathbb{P}_{0}$ for all $t\in\mathbb{R}$ (see also Def. 9.7.(v) in [klenkeProbabilityTheoryComprehensive2013]). The following proposition provides a corresponding mathematical statement, which relies on known results used in the proof of Stone’s theorem and is only stated for the convenience of the reader.
Proposition 1
Let $\hat{H}$, $\Psi_{0}$, and $(\mathbb{P}_{t})_{t\in\mathbb{R}}$ satisfy the assumptions of Thm. 1 above.
  1. 1)
    Then for any $t\in\mathbb{R}$ there exists an element $\partial\Psi_{t}/\partial t$ in $L^{2}(\mathbb{R}^{3N},\mathbb{C})$ such that
    $\lim_{\varepsilon\to 0}\left\|\frac{\partial\Psi_{t}}{\partial t}-\frac{\Psi_{t+\varepsilon}-\Psi_{t}}{\varepsilon}\right\|=0\ .$ (6.5a)
    The map $\Psi\colon t\mapsto\Psi_{t}$ satisfies Eq. (5.3).
  2. 2)
    If, in addition, the family of probability measures $(\mathbb{P}_{t})_{t\in\mathbb{R}}$ is stationary, then there exists a constant $\mathcal{E}\in\mathbb{R}$ such that
    $\hat{H}\Psi_{0}=\mathcal{E}\Psi_{0}\ .$ (6.5b)
    Moreover, $E$, as given via Eq. (6.3c), satisfies $E_{t}=\mathcal{E}$ for all $t\in\mathbb{R}$.
$\diamondsuit$
Again, the proof is given in the appendix.
@@
We consider the simplest quantum-mechanical model of the hydrogen atom: $N=1$ and $V(\vec{r})=-\text{e}^{2}/(4\pi\varepsilon_{0}\left\lvert\vec{r}\right\rvert)$. Using physics conventions, we denote the electron mass by $\mu$ and each stationary basis (wave) function by $\Psi_{nlm}$; $n$, $l$, and $m$ are the usual, appropriately chosen integers and $\mathcal{E}_{n}$ is the energy eigenvalue for the given $n$ (cf. Eq. (6.5b)).
We indeed have $\Psi_{nlm}\in\operatorname{dom}\hat{H}=H^{2}(\mathbb{R}^{3},\mathbb{C})$. Recalling that $\Psi_{nlm}(t,\vec{r})\equiv e^{-\mathfrak{i}\mathcal{E}_{n}t/\hbar}\,\Psi_{nlm}(\vec{r})$, the respective random variables from Thm. 1.2 are given via
@@
It is worth noting that the ‘observables’ $\vec{v}_{a}$ and $E$ from Thm. 1 depend on the particular ensemble one considers (as encoded by $\Psi_{t}$). This is in stark contrast to the momentum and energy operator in quantum mechanics, which generally only depend on the model, not the particular ensemble.
Finally, to physically compare geometric quantum theory and quantum mechanics, one needs to consider cases for which the theories make different predictions. The following example shows that such cases exist.
Example 3
Continuing Ex. 3 above, we consider the initial wave function
$\Psi_{0}=a_{1}\,\Psi_{n_{1}l_{1}m_{1}}+a_{2}\,\Psi_{n_{2}l_{2}m_{2}}$ (6.7a)
with $a_{1}$, $a_{2}\in\mathbb{C}$ satisfying
$\left\lvert a_{1}\right\rvert^{2}+\left\lvert a_{2}\right\rvert^{2}=1\ .$ (6.7b)
For convenience, we denote the first summand of Eq. (6.7a) by $\xi_{1}$, the second one by $\xi_{2}$, and set $\omega=\mathcal{E}_{n_{2}}-\mathcal{E}_{n_{1}}$.
We now determine the probability distribution for the energy in both theories, as given by Tab. 2.
In quantum mechanics, the probability to “measure” $\mathcal{E}_{n_{1}}$ is $\left\lvert a_{1}\right\rvert^{2}$, for $\mathcal{E}_{n_{2}}$ it is $\left\lvert a_{2}\right\rvert^{2}$, and for any other energy it is zero. Hence, for $n_{1}\neq n_{2}$ and nonzero $a_{1}$ and $a_{2}$ the distribution is supported on the set $\{\mathcal{E}_{n_{1}},\mathcal{E}_{n_{2}}\}$ and it is independent of $t$.
In geometric quantum theory, on the other hand, for each $t\in\mathbb{R}$ the energy is given by the random variable
[…]
and the distribution is $\mathbb{P}_{t}\circ(E_{t})^{-1}$. If $n_{1}=n_{2}$ or if either $a_{1}$ or $a_{2}$ are zero, the quantum-mechanical prediction is reproduced. Yet otherwise, there is a qualitative difference, for the distribution depends on $t$ and, generally, energies between $\mathcal{E}_{n_{1}}$ and $\mathcal{E}_{n_{2}}$ may be attained. $\diamondsuit$
It is important to note that it is not possible to put the differing predictions in Ex. 3 above to experimental scrutiny, for the model describes an isolated hydrogen atom (a “bound state”). Generally speaking and as discussed in more detail in part III of this series, to predict the outcomes of “measurements” other particles or radiation probing the system need to be included in the model (see Sec. 5.7 in [sakuraiModernQuantumMechanics1985] for an example). The model of Ex. 2 may be used for making predictions in this sense, if one employs the electron to ‘probe’ the proton instead. That is, one may be able to obtain differing empirical predictions by considering time-dependent electron-proton scattering instead.